# Tsallis relative $\alpha$ entropy of coherence dynamics in Grover’s search algorithm

### 15 Apr 2026 v1

@@ Abstract
Quantum coherence plays a central role in Grover’s search algorithm. We study the Tsallis relative $\alpha$ entropy of coherence dynamics of the evolved state in Grover’s search algorithm. We prove that the Tsallis relative $\alpha$ entropy of coherence decreases with the increase of the success probability, and derive the complementarity relations between the coherence and the success probability. We show that the operator coherence of the first $H^{\otimes n}$ relies on the size of the database NN, the success probability and the target states. Moreover, we illustrate the relationships between coherence and entanglement of the superposition state of targets, as well as the production and deletion of coherence in Grover iterations.
@@ 2 Tsallis relative $\alpha$ entropy of coherence in GSA and complementarity relations
In Section 2.1, we first review the concepts of the Tsallis relative $\alpha$ entropy of coherence and GSA, and then study the coherence dynamics in GSA. We investigate the complementarity relations between coherence and success probability in Section $2.$
@@ 2.1 Tsallis relative $\alpha$ entropy of coherence in GSA
Here $A\simeq B$ means that AA asymptotically equals to $B$ under the condition $t\ll N$. When the success probability is increasing, according to Eqs. (7) and (8), we obtain that the coherence is decreasing, i.e., coherence decreases with the increase of the success probability. In particular, from Eq. (6), when the success probability reaches 1, the coherence is $C_{\alpha}(\rho_{k})=\frac{1}{\alpha-1}(t^{1-\frac{1}{\alpha}}-1)$. In other words, the success probability relies on the coherence depletion in terms of Tsallis relative $\alpha$ entropy of coherence.
@@
According to Eqs. (12) and (13), when $\alpha\in(0,1)$ and $P_{k}=1$, normalization of coherence satisfies $N(C_{\alpha})+t^{1-\frac{1}{\alpha}}\simeq 1$. When $\alpha\in(1,2]$ and $P_{k}=1$, normalization of coherence satisfies $N(C_{\alpha})\simeq 0$. Noting that $N(C_{\alpha})$ is the normalization of $C_{\alpha}$, and $A\simeq B$ means that AA asymptotically equals to $B$ under the condition $t\ll N$, this does not necessarily mean that $C_{\alpha}$ is asymptotically equals to 0, let alone it is equal to 0.
@@ 3 Dynamics of the Tsallis relative $\alpha$ entropy of coherence in GSA
In the computational basis, the NN-dimensional Hadamard matrix has the following three properties: (i) The common coefficient is $\frac{1}{\sqrt{N}}$. (ii) The elements of the first row and the first column are 1. (iii) In any other rows or columns, a half of the elements are 1, and the others are 1-1. Let $H_{y,x}$ denote the element of the $y$th row and the $x$th column in a Hadamard matrix and $t_{y}$ denote the number of $H_{y,x}=1$, $t_{y}=|\{H_{y,x}|H_{y,x}=1,x\in{\{x_{s}\}}\}|$. Denote by $|\psi_{kH_{O}}\rangle$ the state after $H_{O}$ is applied on $|\psi_{kO}\rangle$ and $|\psi_{kH_{P}}\rangle$ the state after $H_{P}$ is applied on $PH^{\otimes n}|\psi_{kO}\rangle$. We have
@@
respectively. Based on Theorem 2, it is easy to see that the coherence of the state $|\psi_{kH_{O}}\rangle$ is related to the size of the database NN, success probability $P_{k}$ and target states. The coherence of the state $|\psi_{kH_{P}}\rangle$ is dependent on NN and $P_{k}$.
@@ 4 Different target states
Conjecture The Tsallis relative $\alpha$ entropy of coherence of the state $|\psi_{kH_{O}}\rangle$ relies on the size of the database NN, the success probability and the target states.
@@ 5 Production and depletion of Tsallis relative $\alpha$ entropy of coherence
According to Theorem 4, the operator coherence between two consecutive iterations of $H_{P}$ and $G$ is depleted. Both of them rely on the size of the database NN and the success probability. Correspondingly, the operator coherence between two consecutive iterations of $H_{O}$ is produced, which relies on the size of the database NN, the success probability and target states.
@@ 6 Comparison with previous works
In this work, we study coherence dynamics in GSA based on Tsallis relative $\alpha$ entropy. The amount of coherence of first $H^{\otimes n}$ depends on the size of the database NN, success probability and target states, and the coherence of two $H^{\otimes n}$ have different effects that one depletes coherence and the other produces coherence. Coherence is not always produced or depleted, but depleted and produced in turn. When $\alpha\in(0,1)$, the coherence is smaller when the superposition state of targets is an entangled one, and the coherence reduction is not monotonic, but related to the parameter $\alpha$ with the increase of the success probability, which are different from [69].
@@ 7 Conclusions and discussions
We have explored the coherence dynamics in Grover’s search algorithm (GSA) based on Tsallis relative $\alpha$ entropy for $\alpha\in(0,1)\cup(1,2]$, and proved that the coherence decreases with the increase of the success probability. We have derived the complementarity relations between Tsallis relative $\alpha$ entropy of coherence and the success probability. Moreover, we have studied how each basic operator contributes to the coherence in GSA, and proved that the amount of operator coherence of $H_{O}$ relies on the size of the database NN, the success probability and the target states. Following the idea in [69], we have also discussed the operator coherence of $|\psi_{kH_{O}}\rangle$ for different target states. Finally, when $\alpha\in(1,2]$, we have derived the variations of operator coherence between two consecutive iterations of $H_{O}$, $H_{P}$ and $G$ in GSA, and the variations of suboperator coherence of each basic operator $H_{O}$ and $H_{P}$ in one Grover iteration. The operators $H_{O}$ and $H_{P}$ have different effects on coherence, one produces coherence and another depletes it. Coherence of the $H^{\otimes n}$ is not always depleted, but depleted and produced alternatively. It oscillates during Grover’s search algorithm application.

### 16 Apr 2026 v2

@@ Abstract
Quantum coherence plays a central role in Grover’s search algorithm. We study the Tsallis relative $\alpha$ entropy of coherence dynamics of the evolved state in Grover’s search algorithm. We prove that the Tsallis relative $\alpha$ entropy of coherence decreases with the increase of the success probability, and derive the complementarity relations between the coherence and the success probability. We show that the operator coherence of the first $H^{\otimes n}$ relies on the size of the database $N$, the success probability and the target states. Moreover, we illustrate the relationships between coherence and entanglement of the superposition state of targets, as well as the production and deletion of coherence in Grover iterations.
@@ 2 Tsallis relative $\alpha$ entropy of coherence in GSA and complementarity relations
In Section 2.1, we first review the concepts of the Tsallis relative $\alpha$ entropy of coherence and GSA, and then study the coherence dynamics in GSA. We investigate the complementarity relations between coherence and success probability in Section 2.2.
@@ 2.1 Tsallis relative $\alpha$ entropy of coherence in GSA
Here $A\simeq B$ means that $A$ asymptotically equals to $B$ under the condition $t\ll N$. When the success probability is increasing, according to Eqs. (7) and (8), we obtain that the coherence is decreasing, i.e., coherence decreases with the increase of the success probability. In particular, from Eq. (6), when the success probability reaches 1, the coherence is $C_{\alpha}(\rho_{k})=\frac{1}{\alpha-1}(t^{1-\frac{1}{\alpha}}-1)$. In other words, the success probability relies on the coherence depletion in terms of Tsallis relative $\alpha$ entropy of coherence.
@@
According to Eqs. (12) and (13), when $\alpha\in(0,1)$ and $P_{k}=1$, normalization of coherence satisfies $N(C_{\alpha})+t^{1-\frac{1}{\alpha}}\simeq 1$. When $\alpha\in(1,2]$ and $P_{k}=1$, normalization of coherence satisfies $N(C_{\alpha})\simeq 0$. Noting that $N(C_{\alpha})$ is the normalization of $C_{\alpha}$, and $A\simeq B$ means that $A$ asymptotically equals to $B$ under the condition $t\ll N$, this does not necessarily mean that $C_{\alpha}$ is asymptotically equals to 0, let alone it is equal to 0.
@@ 3 Dynamics of the Tsallis relative $\alpha$ entropy of coherence in GSA
In the computational basis, the $N$-dimensional Hadamard matrix has the following three properties: (i) The common coefficient is $\frac{1}{\sqrt{N}}$. (ii) The elements of the first row and the first column are 1. (iii) In any other rows or columns, a half of the elements are 1, and the others are 1-1. Let $H_{y,x}$ denote the element of the $y$th row and the $x$th column in a Hadamard matrix and $t_{y}$ denote the number of $H_{y,x}=1$, $t_{y}=|\{H_{y,x}|H_{y,x}=1,x\in{\{x_{s}\}}\}|$. Denote by $|\psi_{kH_{O}}\rangle$ the state after $H_{O}$ is applied on $|\psi_{kO}\rangle$ and $|\psi_{kH_{P}}\rangle$ the state after $H_{P}$ is applied on $PH^{\otimes n}|\psi_{kO}\rangle$. We have
@@
respectively. Based on Theorem 2, it is easy to see that the coherence of the state $|\psi_{kH_{O}}\rangle$ is related to the size of the database $N$, success probability $P_{k}$ and target states. The coherence of the state $|\psi_{kH_{P}}\rangle$ is dependent on $N$ and $P_{k}$.
@@ 4 Different target states
Conjecture The Tsallis relative $\alpha$ entropy of coherence of the state $|\psi_{kH_{O}}\rangle$ relies on the size of the database $N$, the success probability and the target states.
@@ 5 Production and depletion of Tsallis relative $\alpha$ entropy of coherence
According to Theorem 4, the operator coherence between two consecutive iterations of $H_{P}$ and $G$ is depleted. Both of them rely on the size of the database $N$ and the success probability. Correspondingly, the operator coherence between two consecutive iterations of $H_{O}$ is produced, which relies on the size of the database $N$, the success probability and target states.
@@ 6 Comparison with previous works
In this work, we study coherence dynamics in GSA based on Tsallis relative $\alpha$ entropy. The amount of coherence of first $H^{\otimes n}$ depends on the size of the database $N$, success probability and target states, and the coherence of two $H^{\otimes n}$ have different effects that one depletes coherence and the other produces coherence. Coherence is not always produced or depleted, but depleted and produced in turn. When $\alpha\in(0,1)$, the coherence is smaller when the superposition state of targets is an entangled one, and the coherence reduction is not monotonic, but related to the parameter $\alpha$ with the increase of the success probability, which are different from [69].
@@ 7 Conclusions and discussions
We have explored the coherence dynamics in Grover’s search algorithm (GSA) based on Tsallis relative $\alpha$ entropy for $\alpha\in(0,1)\cup(1,2]$, and proved that the coherence decreases with the increase of the success probability. We have derived the complementarity relations between Tsallis relative $\alpha$ entropy of coherence and the success probability. Moreover, we have studied how each basic operator contributes to the coherence in GSA, and proved that the amount of operator coherence of $H_{O}$ relies on the size of the database $N$, the success probability and the target states. Following the idea in [69], we have also discussed the operator coherence of $|\psi_{kH_{O}}\rangle$ for different target states. Finally, when $\alpha\in(1,2]$, we have derived the variations of operator coherence between two consecutive iterations of $H_{O}$, $H_{P}$ and $G$ in GSA, and the variations of suboperator coherence of each basic operator $H_{O}$ and $H_{P}$ in one Grover iteration. The operators $H_{O}$ and $H_{P}$ have different effects on coherence, one produces coherence and another depletes it. Coherence of the $H^{\otimes n}$ is not always depleted, but depleted and produced alternatively. It oscillates during Grover’s search algorithm application.